THE ENERGY LEVELS OF URANIUM IV, V, AND VI

\underline{SCOTT R. BROZELL}, PHILLIP A. CHRISTIANSEN, and RUSSELL M. PITZER, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Spin-orbit configuration-interaction calculations are performed on U^{3+}, U^{4+}, and U^{5+}. Uranium is modeled with relativistic effective core potentials of various sizes and Gaussian cc-pVDZ and cc-pVTZ basis sets. The U^{4+} 5f^{2} ion has been studied extensively, and the calculations are compared with the experimental spectrum^{a}.

aJ. F. Wyart, V Kaufman, and J. Sugar Phys. Scr. 22, 389 (1980)